Panamint Systems Corporation is estimating activity costs associated with producing disk drives, tapes drives, and wire drives. The indirect labor can be traced to four separate activity pools The budgeted activity cost and activity base data by product are provided below Activity Cost Activity Base Number of purchase orders Number of production orders Number of moves Procurement333,500 Scheduling $245,900 Materials handling Product development Production ,513,200 $424,700 $743,700 Number of engineering changes Machine hours Number Number of Number Number of Number of Units of Production of Engineering Machine Purchase Orders Moves Changes Hours Orders isK 4,020 320 235 780 4,300 20 11200 2,800 1,400 14 1,700 1,700 drives Tape 1,900 500 8,000 3,900 drives Wire drives 11,200 Determine the activity rate for production per machine hour

Oa. $19.48 Ob. $140.9
Answer #1

Activity base for production is machine hours.

Total machine hours = 1,700 + 8,000 + 11,200 = 20,900

Activity rate For production = Total activity cost of production / Total machine hours = $1,513,200 / 20,900 machine hours = $72.4 [ option c ]